# references.bib

### v2

@@ I Introduction
WiFi channel state information is typically computed in the process of equalizing received symbols using known sync words and pilot carriers. However, especially in multi-carrier systems, it contains rich information about the propagation environment. The ubiquity of WiFi therefore makes CSI an attractive source of data for sensing applications. Data-driven methods for recovering physical information from the CSI of a multi-carrier wireless link are well studied. With the power of machine learning techniques, researchers have succeeded in accomplishing a wide variety of sensing tasks, including pose estimation and activity classification [ma_wifi_2019], and with sufficient bandwidth, even as finely-grained tasks as sensing human breathing at a distance. While existing research has confirmed the feasibility, and indeed the surprising practicality, of using CSI to predict low-dimensional indicators of the physical world, work showing high-dimensional environment reconstruction capability is comparatively scarce.
@@
Given these challenges, many prior works have adopted generative adversarial networks (GANs) [goodfellow_generative_2014] as a solution for CSI-based image generation. The GAN framework trains two models against each other: a generator that learns the target distribution and a discriminator that learns to distinguish generated samples from real samples. GAN-based approaches were some of the first to be applied to generating images from CSI; one of the first direct CSI to image works [kato_csi2image_2021] developed a GAN to generate a 64x64 RGB image. The authors demonstrated their method correctly classifying objects and positioning human subjects, but they noted poor performance on humans in motion. Subsequent work in [yu_wifi-based_2022] and [yu_rfgan_2023] applies the GAN framework to transform images using information determined from CSI using pose keypoints determined from CSI, which circumvents the challenge of generating images from scratch. While GANs have shown promise, they introduce significant complexity into the training process, as they require training two models simultaneously and careful tuning to ensure stability and convergence [arjovsky_towards_2017]. This complexity motivates the exploration of simpler models with more straightforward training procedures. A recent non-GAN approach addressing this issue is presented in [strohmayer_through-wall_2024], where the authors propose a mixture-of-product-of-experts variational autoencoder (VAE) that jointly trains an MLP-based CSI encoder and a ground-truth image encoder.
@@
To address these limitations, we propose LatentCSI, a novel image generation scheme that leverages a pretrained latent diffusion model (LDM). Diffusion models have recently gained attention for their ability to produce high-quality images through iterative denoising guided by text embeddings. Latent diffusion models in particular operate in a lower-dimensional latent space rather than the color pixel space, enabling faster and more memory-efficient image synthesis.
Compared to existing methods, LatentCSI adopts a simpler design that trains a single, lightweight neural network to predict latent embeddings directly from CSI data. The reduced dimensionality of the output, combined with the expressive power of a pretrained LDM diffusion model and decoder allows for the generation of high-resolution images at a lower computational cost. Moreover, since the LDM supports text-guided image generation, our approach naturally enables an optional refinement step, in which the generated image can be modified according to textual instructions—for example, altering a person’s face or clothing appearance to match a given description. Importantly, the low-dimensional bottleneck in our encoding-decoding pipeline inherently reduces the risk of reproducing sensitive fine-grained details such as facial features, clothing, or background elements. At the same time, the LDM can synthesize visually plausible content aligned with user intent via text-guided modification, striking a balance between interpretability and privacy. This capability enables controllable and privacy-preserving image generation while preserving a meaningful correspondence with the underlying CSI signals.
The primary contribution of this work is the introduction of an LDM-based framework for generating images of the physical environment from CSI data. Our approach achieves state-of-the-art resolution and text-conditioned image synthesis with a simple architecture, a straightforward training procedure, and low computational overhead. We validate our method using two paired CSI–image datasets: one collected through experiments involving variations in a walking subject’s position and orientation; and another based on a subset of the publicly available MM-Fi dataset [yang_mm-fi_2023], selected to represent diverse human poses with less visual complexity. For both datasets, we qualitatively evaluate the generated images and quantitatively compare our approach against a baseline model trained end-to-end on full-resolution images.
@@ II-A System model
We consider a WiFi sensing system designed to monitor indoor environments by capturing human activity through wireless signals. The system consists of a WiFi access point, a WiFi terminal, and an RGB camera that is available only during the training phase. This setup enables the collection of paired channel state information (CSI) and image data, which are used to train an image generation model. LatentCSI, is intended for practical scenarios in which camera imagery can be collected temporarily—such as during initial system calibration—but becomes unavailable during deployment due to obstructions or privacy concerns. As a result, while the model leverages visual data during training, inference is performed exclusively using CSI, making the approach suitable for privacy-sensitive or camera-limited environments.
@@
II-B Overview of LatentCSI
Our proposed method, LatentCSI, builds upon a pretrained latent diffusion model, Stable Diffusion v1.5 [rombach_high-resolution_2022]. To clarify our design, we first describe the standard inference process of LDMs. We proceed to introduce our modified inference procedure, which adapts an image-to-image pipeline to use CSI instead of images. The training procedure of the key component in this pipeline—the CSI encoder that predicts the LDM’s latent representation—is presented in the following section.
@@ II-B1 LDM image-to-image pipeline
Diffusion models generate images by reversing a diffusion process, progressively transforming sampled Gaussian noise into an image through iterative denoising guided by conditioning inputs such as text prompts. In contrast, the image-to-image procedure, popularized in [meng_sdedit_2021], starts from an input image, adds noise to simulate forward diffusion up to a chosen intermediate timestep, and then denoises to produce the output.
Latent diffusion models perform denoising in latent space rather than pixel space, offering performance benefits. Since our implementation of LatentCSI is based on Stable Diffusion, we describe the LDM’s operation according to the implementation of Stable Diffusion. While different implementations of LDM may vary in model architecture, their fundamental operation remains the same. The image-to-image pipeline of an LDM is illustrated in Figure 1.
@@
II-B2 LatentCSI Inference Process
Figure 2 shows the inference pipeline of LatentCSI. The key difference from the standard LDM pipeline is that the VAE encoder is replaced with a CSI encoder, enabling the LDM’s latent starting point to be estimated directly from CSI instead of an image. Similar to an LDM, data in the LatentCSI pipeline is processed primarily in latent space, in contrast to existing CSI-to-image generation methods that directly map CSI to image space. This latent-space design reduces computational cost, as the latent embedding has much lower dimensionality than high-resolution images.
@@ II-B2 LatentCSI Inference Process
Figure 2: LatentCSI: img2img with custom CSI encoder
II-C LatentCSI training process
@@
The overall training procedure is illustrated in Figure 3. The CSI encoder learns to regress directly to the mean latent representation produced by the pretrained VAE encoder for each corresponding image $y$. In our approach, the CSI input $x$ uses only amplitude information. Given a complex-valued CSI measurement $x_{c}\in\mathbb{C}^{s}$, the input is computed as $x=\sqrt{\text{Re}(x_{c})^{2}+\text{Im}(x_{c})^{2}}$. This choice follows prior work [ma_wifi_2019] [yu_wifi-based_2022] [strohmayer_through-wall_2024] suggesting that amplitude alone is sufficient for many sensing tasks, and avoids the additional complexity of handling phases in unsynchronized device configurations.
@@ II-D Model architecture
We first apply a fully connected layer to map the 1D CSI input of size $s$ to a a 1D vector before reshaping it into a tensor with $b$ channels, where $b$ is a tunable parameter. We use 4 upsampling steps that each halves the number of channels and doubles the resolution. We then use a final convolutional layer to reduce the number of channels to the output specification. Each upsampling layer consists of two residual blocks and a transpose convolutional layer. Each residual block is itself composed of two convolutional layers. The last 3 upsampling steps use a cross attention layer to ensure that later spatially upsampling blocks attend to global information. This architecture strongly resembles that of the Stable Diffusion VAE encoder, which uses 4 similar downsampling steps followed by convolutional layers.
@@ III Evaluation
We evaluate LatentCSI on two datasets. For each dataset, we present a qualitative analysis using sample reference and generated images from the test sets withheld from the model.
@@
For both datasets, we report these metrics in two cases: (1) using the entire generated and reference images, and (2) using cropped image regions containing human figures, as identified by the YOLOv3 object detection model [redmon_yolov3_2018]. Given the human-centric nature of the datasets, the second case provides a focused evaluation of the model’s ability to preserve semantically important content. In both evaluation schemes, we benchmark LatentCSI against a baseline model with the same architecture and similar parameter count, which directly predicts RGB pixel values from CSI without the use of latent mechanisms.
Both LatentCSI and the baseline model were trained using an early stopping criterion for validation loss stagnation over 5 consecutive training epochs. We used the ADAM optimizer with a fixed learning rate of 0.0005. To account for variability arising from stochastic weight initialization and training dynamics, each model was trained 5 times. Image samples are generated using the model instance with the best test loss. All experiments were conducted on a system equipped with a single NVIDIA H100 GPU and eight CPU cores of an AMD EPYC 9654 processor.
@@ III-A Dataset 1: Position & orientation of one moving human subject
We first evaluate the proposed method on a custom dataset designed to assess performance in predicting the position and orientation of a moving human subject. Data was collected from a single individual walking back and forth in a small indoor environment. CSI measurements were recorded at 10 Hz over 25 minutes using two ASUS NUCs equipped with Intel AX210 wireless cards, operating over a 160 MHz channel with a single transmit-receive antenna pair. A total of 1992 subcarriers were captured per sample using the FeitCSI tool [feitcsi:project]. Each CSI sample was paired with a synchronized $640\times 480$ RGB image captured via an Intel Realsense D435 camera. The resulting 15,000 image-CSI pairs were post-processed by cropping and resampling images to $512\times 512$ resolution. Facial features were blurred only in the presented examples; unaltered images were used for training and inference. The dataset was split into 80% training, 10% validation, and 10% testing. The CSI encoder model was configured with $b=256$ initial channels, yielding $22{,}914{,}052$ trainable parameters. The baseline model used $b=8$ and comprised $16{,}434{,}395$ parameters.
@@
Reference Proposed Baseline
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table I presents representative samples from the test set, while quantitative results are shown in Table III. LatentCSI exhibits marginal improvements over the baseline in RMSE and SSIM, but achieves a substantially lower FID, indicating superior perceptual quality. This advantage is further amplified when evaluation is restricted to regions containing human figures, suggesting better reconstruction of salient information. Despite comparable model complexity, the baseline model—which directly predicts pixel values—requires approximately three times longer to train, owing to the increased output dimensionality. Dataset 1 presents a realistic, unsanitized setting with diverse views of a moving subject and frequent occlusions. These factors increase pose significant challenges for pixel-space prediction.
Qualitative results indicate that the baseline model underfits the dataset, likely constrained by its limited capacity to model high-resolution pixel distributions. In contrast, LatentCSI produces better reconstructions. These outputs serve as effective initialization points for downstream diffusion-based refinement, as shown in Table IV.
@@ III-B Dataset 2: Human poses with MM-Fi
We evaluate our proposed method on the popular MM-Fi dataset [yang_mm-fi_2023]. Data is collected from 3 antennas each using 114-carrier 40 MHz channels, resulting in 342 total subcarriers. Compared to Dataset 1, this offers lower spectral resolution but greater spatial diversity. The selected subset comprises environment 3, the first two subjects, and four arm movement activities: 13–14 (raising hand, left/right) and 17–18 (waving hand, left/right), totaling 23,760 samples. The proposed model is configured with $b=256$ and has 13,621,252 parameters; the baseline uses $b=32$ with 11,490,275 parameters.
@@
Reference Proposed Baseline
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
However, the proposed method outperforms the baseline in FID, indicating superior perceptual quality. When evaluation is restricted to regions containing human subjects, the RMSE and SSIM differences narrow, while the FID advantage of the proposed model approximately doubles. Qualitative analysis suggests the baseline model tends to produce artifacts and anatomically implausible outputs. In contrast, the latent-space model yields more coherent representations of salient features.
@@
[…]

### v3

@@ I Introduction
WiFi channel state information is typically computed in the process of equalizing received symbols using known sync words and pilot carriers. However, especially in multi-carrier systems, it contains rich information about the propagation environment. The ubiquity of WiFi therefore makes CSI an attractive source of data for sensing applications. Data-driven methods for recovering physical information from the CSI of a multi-carrier wireless link are well studied. With the power of machine learning techniques, researchers have succeeded in accomplishing a wide variety of sensing tasks, including pose estimation and activity classification [1], and with sufficient bandwidth, even as finely-grained tasks as sensing human breathing at a distance. While existing research has confirmed the feasibility, and indeed the surprising practicality, of using CSI to predict low-dimensional indicators of the physical world, work showing high-dimensional environment reconstruction capability is comparatively scarce.
@@
Given these challenges, many prior works have adopted generative adversarial networks (GANs) [2] as a solution for CSI-based image generation. The GAN framework trains two models against each other: a generator that learns the target distribution and a discriminator that learns to distinguish generated samples from real samples. GAN-based approaches were some of the first to be applied to generating images from CSI; one of the first direct CSI to image works [3] developed a GAN to generate a 64x64 RGB image. The authors demonstrated their method correctly classifying objects and positioning human subjects, but they noted poor performance on humans in motion. Subsequent work in [4] and [5] applies the GAN framework to transform images using information determined from CSI using pose keypoints determined from CSI, which circumvents the challenge of generating images from scratch. While GANs have shown promise, they introduce significant complexity into the training process, as they require training two models simultaneously and careful tuning to ensure stability and convergence [6]. This complexity motivates the exploration of simpler models with more straightforward training procedures. A recent non-GAN approach addressing this issue is presented in [7], where the authors propose a mixture-of-product-of-experts variational autoencoder (VAE) that jointly trains an MLP-based CSI encoder and a ground-truth image encoder.
@@
To address these limitations, we propose novel image generation scheme that leverages a pretrained latent diffusion model (LDM). Diffusion models have recently gained attention for their ability to produce high-quality images through iterative denoising guided by text embeddings. Latent diffusion models in particular operate in a lower-dimensional latent space rather than the color pixel space, enabling faster and more memory-efficient image synthesis.
Compared to existing methods, opts a simpler design that trains a single, lightweight neural network to predict latent embeddings directly from CSI data. The reduced dimensionality of the output, combined with the expressive power of a pretrained LDM diffusion model and decoder allows for the generation of high-resolution images at a lower computational cost. Moreover, since the LDM supports text-guided image generation, our approach naturally enables an optional refinement step, in which the generated image can be modified according to textual instructions—for example, altering a person’s face or clothing appearance to match a given description. Importantly, the low-dimensional bottleneck in our encoding-decoding pipeline inherently reduces the risk of reproducing sensitive fine-grained details such as facial features, clothing, or background elements. At the same time, the LDM can synthesize visually plausible content aligned with user intent via text-guided modification, striking a balance between interpretability and privacy. This capability enables controllable and privacy-preserving image generation while preserving a meaningful correspondence with the underlying CSI signals.
The primary contribution of this work is the introduction of an LDM-based framework for generating images of the physical environment from CSI data. Our approach achieves state-of-the-art resolution and text-conditioned image synthesis with a simple architecture, a straightforward training procedure, and low computational overhead. We validate our method using two paired CSI–image datasets: one collected through experiments involving variations in a walking subject’s position and orientation; and another based on a subset of the publicly available MM-Fi dataset [8], selected to represent diverse human poses with less visual complexity. For both datasets, we qualitatively evaluate the generated images and quantitatively compare our approach against a baseline model trained end-to-end on full-resolution images and the hybrid GAN strategy introduced in CSI2Image [3].
@@ II-A System model
We consider a WiFi sensing system designed to monitor indoor environments by capturing human activity through wireless signals. The system consists of a WiFi access point, a WiFi terminal, and an RGB camera that is available only during the training phase. This setup enables the collection of paired channel state information (CSI) and image data, which are used to train an image generation model. s intended for practical scenarios in which camera imagery can be collected temporarily—such as during initial system calibration—but becomes unavailable during deployment due to obstructions or privacy concerns. As a result, while the model leverages visual data during training, inference is performed exclusively using CSI, making the approach suitable for privacy-sensitive or camera-limited environments.
@@
II-B Overview of missing
Our proposed method, uilds upon a pretrained latent diffusion model, Stable Diffusion v1.5 [9]. To clarify our design, we first describe the standard inference process of LDMs. We proceed to introduce our modified inference procedure, which adapts an image-to-image pipeline to use CSI instead of images. The training procedure of the key component in this pipeline—the CSI encoder that predicts the LDM’s latent representation—is presented in the following section.
@@ II-B1 LDM image-to-image pipeline
Diffusion models generate images by reversing a diffusion process, progressively transforming sampled Gaussian noise into an image through iterative denoising guided by conditioning inputs such as text prompts. In contrast, the image-to-image procedure, popularized in [10], starts from an input image, adds noise to simulate forward diffusion up to a chosen intermediate timestep, and then denoises to produce the output.
Latent diffusion models perform denoising in latent space rather than pixel space, offering performance benefits. Since our implementation of based on Stable Diffusion, we describe the LDM’s operation according to the implementation of Stable Diffusion. While different implementations of LDM may vary in model architecture, their fundamental operation remains the same. The image-to-image pipeline of an LDM is illustrated in Figure 1.
@@
II-B2 ference Process
Figure 2 shows the inference pipeline of he key difference from the standard LDM pipeline is that the VAE encoder is replaced with a CSI encoder, enabling the LDM’s latent starting point to be estimated directly from CSI instead of an image. Similar to an LDM, data in the peline is processed primarily in latent space, in contrast to existing CSI-to-image generation methods that directly map CSI to image space. This latent-space design reduces computational cost, as the latent embedding has much lower dimensionality than high-resolution images.
@@
Figure 2: mg2img with custom CSI encoder
II-C aining process
@@
The overall training procedure is illustrated in Figure 3. The CSI encoder learns to regress directly to the mean latent representation produced by the pretrained VAE encoder for each corresponding image $y$. In our approach, the CSI input $x$ uses only amplitude information. Given a complex-valued CSI measurement $x_{c}\in\mathbb{C}^{s}$, the input is computed as $x=\sqrt{\text{Re}(x_{c})^{2}+\text{Im}(x_{c})^{2}}$. This choice follows prior work [1] [4] [7] suggesting that amplitude alone is sufficient for many sensing tasks, and avoids the additional complexity of handling phases in unsynchronized device configurations.
@@ II-D Model architecture
We first apply a fully connected layer to map the 1D CSI input of size $s$ to a 1D vector before reshaping it into a tensor with $b$ channels, where $b$ is a tunable parameter. We use 4 upsampling steps that each halves the number of channels and doubles the resolution. We then use a final convolutional layer to reduce the number of channels to the output specification. Each upsampling layer consists of two residual blocks and a transpose convolutional layer. Each residual block is itself composed of two convolutional layers. The last 3 upsampling steps use a cross attention layer to ensure that later spatially upsampling blocks attend to global information. This architecture strongly resembles that of the Stable Diffusion VAE encoder, which uses 4 similar downsampling steps followed by convolutional layers.
@@ III Evaluation
We evaluate two datasets. For each dataset, we present a qualitative analysis using sample reference and generated images from the test sets withheld from the model.
@@
For both datasets, we report these metrics in two cases: (1) using the entire generated and reference images, and (2) using cropped image regions containing human figures, as identified by the YOLOv3 object detection model [11]. Given the human-centric nature of the datasets, the second case provides a focused evaluation of the model’s ability to preserve semantically important content.
In both evaluation schemes, we benchmark ainst two strategies: a baseline model using the exact same architecture and training procedure as LatentCSI but predicting pixels directly; and the “hybrid” GAN strategy (K=8) from CSI2Image[3] which combines supervised loss with adversarial conditional GAN loss.
d the baseline model were trained using an early stopping criterion for validation loss stagnation over 5 consecutive training epochs. We used the ADAM optimizer with a fixed learning rate of 0.0005. To account for variability arising from stochastic weight initialization and training dynamics, each model was trained 5 times. Image samples are generated using the model instance with the best test loss. All experiments were conducted on a system equipped with a single NVIDIA H100 GPU and eight CPU cores of an AMD EPYC 9654 processor.
@@ III-A Dataset 1: Position & orientation of one moving human subject
We first evaluate the proposed method on a custom dataset designed to assess performance in predicting the position and orientation of a moving human subject. Data was collected from a single individual walking back and forth in a small indoor environment. CSI measurements were recorded at 10 Hz over 25 minutes using two ASUS NUCs equipped with Intel AX210 wireless cards, operating over a 160 MHz channel with a single transmit-receive antenna pair. A total of 1992 subcarriers were captured per sample using the FeitCSI tool [12]. Each CSI sample was paired with a synchronized $640\times 480$ RGB image captured via an Intel Realsense D435 camera. The resulting 15,000 image-CSI pairs were post-processed by cropping and resampling images to $512\times 512$ resolution. Facial features were blurred only in the presented examples; unaltered images were used for training and inference. The dataset was split into 80% training, 10% validation, and 10% testing. The CSI encoder model was configured with $b=256$ initial channels, yielding 22,914,05222{,}914{,}052 trainable parameters. The baseline model used $b=8$ and comprised 16,434,39516{,}434{,}395 parameters.
@@
Reference LatentCSI Baseline CSI2Image[3]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table I presents representative samples from the test set, while quantitative results are shown in Table III. hibits marginal improvements over the baseline in RMSE and SSIM, but achieves a substantially lower FID, indicating superior perceptual quality. This advantage is further amplified when evaluation is restricted to regions containing human figures, suggesting better reconstruction of salient information. Despite comparable model complexity, the baseline model—which directly predicts pixel values—requires approximately three times longer to train, owing to the increased output dimensionality. Dataset 1 presents a realistic, unsanitized setting with diverse views of a moving subject and frequent occlusions. These factors increase pose significant challenges for pixel-space prediction.
Qualitative results indicate that the baseline and CSI2Image underfit the dataset, likely constrained by their limited capacity to model high-resolution pixel distributions. In contrast, oduces better reconstructions with a smaller model. These outputs serve as effective initialization points for downstream diffusion-based refinement, as shown in Table IV.
@@ III-B Dataset 2: Human poses with MM-Fi
We evaluate our proposed method on the popular MM-Fi dataset [8]. Data is collected from 3 antennas each using 114-carrier 40 MHz channels, resulting in 342 total subcarriers. Compared to Dataset 1, this offers lower spectral resolution but greater spatial diversity. The selected subset comprises environment 3, the first two subjects, and four arm movement activities: 13–14 (raising hand, left/right) and 17–18 (waving hand, left/right), totaling 23,760 samples. The proposed model is configured with $b=256$ and has 13,621,252 parameters; the baseline uses $b=32$ with 11,490,275 parameters.
@@
Reference LatentCSI Baseline CSI2Image[3]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
However, the proposed method outperforms the baseline in FID, indicating superior perceptual quality. When evaluation is restricted to regions containing human subjects, the RMSE and SSIM differences narrow, while the FID advantage of the proposed model approximately doubles. Qualitative analysis suggests the baseline model tends to produce artifacts and anatomically implausible outputs. The CSI2Image model is significantly underfit, and likely requires architecture tuning to work well with the higher resolution used in this work. In contrast, the latent-space model yields more coherent representations of salient features.
@@
Dataset 1 $134.23\pm 0.07$.18) $18.95\pm 0.00$.44) $87\pm 0.00.00$) $126.09\pm 0.07$.15) $19.81\pm 0.00$.66) $0.83\pm 0.00$.01) $11.9\pm 0.0$.1) $\text{05:02}\pm\text{00:46}$
Dataset 1 Baseline $268.03\pm 0.04$.23) $20.45\pm 0.00$.27) $84\pm 0.00.00$) $296.93\pm 0.04$.83) $21.74\pm 0.00$.29) $78\pm 0.00.00$) $36.4\pm 0.7$.9) $\text{16:29}\pm\text{03:35}$
Dataset 1 CSI2Image $392.46\pm 0.11$.0) $39.14\pm 0.01$.94) $0.58\pm 0.00$.06) $349.87\pm 0.15$.87) $45.57\pm 0.02$.47) $0.49\pm 0.00$.04) $67.8\pm 0.1$.7) $\text{15:28}\pm\text{5:09}$
Dataset 2 $28.21\pm 0.00$.78) $7.87\pm 0.00$.09) $94\pm 0.00.00$) $27.90\pm 0.00$.43) $7.90\pm 0.00$.03) $92\pm 0.00.00$) $17.9\pm 0.0$.2) $\text{16:24}\pm\text{00:03}$
Dataset 2 Baseline $47.67\pm 0.02$.04) $7.15\pm 0.00$.50) $97\pm 0.00.00$) $69.34\pm 0.02$.50) $7.19\pm 0.00$.24) $93\pm 0.00.00$) $59.3\pm 0.0$.2) $\text{91:09}\pm\text{19:19}$
Dataset 2 CSI2Image $312.24\pm 0.21$.5) $24.15\pm 0.00$.64) $0.76\pm 0.00$.03) $372.99\pm 0.10$.25) $49.72\pm 0.00$.65) $0.46\pm 0.00$.02) $73.0\pm 0.2$.3) $\text{32:53}\pm\text{10:09}$